# MetaEarth: A Generative Foundation Model for Global-Scale Remote Sensing Image Generation

### v1

@@ I Introduction
To address the challenge of model capacity, we constructed a world-scale generative foundation model with over 600 million parameters based on the denoising diffusion paradigm. To train the model across diverse spatial resolutions globally, we collected a large-scale remote sensing image dataset comprising images and their geographical information (latitude, longitude, and resolution) from multiple spatial resolutions that cover most regions globally. To the best of our knowledge, this is currently the largest generative model in terms of parameter size in the field of remote sensing.
To generate cross-resolution and controllable remote sensing images, we propose a novel self-cascading generative framework. Different from previous single-stage generation methods [1, 2], the proposed framework generates multi-resolution images from a given geographical location in a recursive manner. In different cascading stages, images are generated sequentially from low to high resolutions. Specifically, we constructed a unified model across different cascading stages, where in each stage, the previously generated low-resolution image, as well as the geographical resolution, are embedded as controllable generative variables to guide the generation of images at higher resolutions. As the cascading stages accumulate, the generated images manifest diversity in both resolution and content, enabling the generation of varied parallel scenarios.
@@ II Related work
II-A Visual Generative Models
@@
II-B Diffusion-Based Image Generation
@@
II-C Remote Sensing Image Generation
@@
III METHODOLOGY
@@
III-A Revisit of Denoising Diffusion Probabilistic Model
@@
III-B The MetaEarth Self-Cascaded Generation Framework
@@
where the image $x_{0}^{(k+m)}$ generated from $(k+m)$-th stage will eventually has a size of $(N^{m}H\times N^{m}W)$ pixels. To achieve the above process with reasonable memory and computational costs, we designed a sliding window generation process and a noise sampling strategy that enables the generation of continuous, unbounded scenes in a memory-efficient manner. These approaches will be detailed in Subsection III-C.
@@
III-C Unbounded Image Generation
@@
III-D Implimentation Details
III-D1 Denoising Network Design
@@
For each feature map $h$ within every ResBlock, we divide the embedding $c_{t}^{(k)}$ into two parts using the chunk function in PyTorch and then scale the feature map $h$ with the chunked embeddings:
$\begin{split}&c_{\text{scale}},c_{\text{shift}}={\texttt{chunk}}(c_{t}^{(k)},2)\\ &h\leftarrow{(1+c_{\text{scale}})}\cdot h+c_{\text{shift}}\end{split}$ (22)
@@ III-D2 Training Details
IV EXPERIMENTS
IV-A Experimental Setup
IV-A1 Dataset
@@ IV-A2 Evaluation Metrics
IV-B Qualitative Analysis
IV-B1 Global-Scale Image Generation
Our experiments show that the proposed MetaEarth can generate a variety of remote sensing scenes worldwide, including glaciers, snowfields, deserts, forests, beaches, farmlands, industrial areas, residential areas, etc. Fig. LABEL:fig:teaser and Fig. 4 showcase some results of our model. Owing to the guidance from the conditional input, the model is capable of generating vivid and diverse images with distinct regional characteristics at different resolutions. We also evaluated the semantical guidance provided by the image conditional input. In Fig. 5, we use large-scale low-resolution images of 256m/pixel in the Americas, Europe, and Asia as image condition guidance, and we can see the model can generate urban images consistent with the regional style.
@@ IV-B1 Global-Scale Image Generation
IV-B2 Multi-Resolution Image Generation
@@
IV-B3 Arbitrary-sized Image Generation
@@
IV-B4 Generalization to Unseen Conditions
@@
IV-B5 Comparison with other text-to-image models
@@ IV-C Ablation Study
IV-C1 Effectiveness of Resolution Guidance
@@
IV-C2 Evaluation on Noise Sampling
@@
IV-D Evaluation on Downstream Tasks

### v2

@@ 1 Introduction
Figure 1: We propose MetaEarth, a generative foundation model that simulates Earth’s visuals from an overhead perspective. MetaEarth shows powerful capabilities on generating worldwide, multi-resolution, unbounded, and virtually limitless remote sensing images. For more visualization and animated results, please refer to our project page.
@@
To address the challenge of model capacity, we constructed a world-scale generative foundation model with over 600 million parameters based on the denoising diffusion paradigm. To train the model across diverse spatial resolutions globally, we collected a large-scale remote sensing image dataset comprising images and their geographical information (latitude, longitude, and resolution) from multiple spatial resolutions that cover most regions globally. To generate cross-resolution and controllable remote sensing images, we propose a novel self-cascading generative framework. Different from previous single-stage generation methods [1, 2], the proposed framework generates multi-resolution images from a given geographical location in a recursive manner. In different cascading stages, images are generated sequentially from low to high resolutions. Specifically, we constructed a unified model across different cascading stages, where in each stage, the previously generated low-resolution image, as well as the geographical resolution, are embedded as controllable generative variables to guide the generation of images at higher resolutions. As the cascading stages accumulate, the generated images manifest diversity in both resolution and content, enabling the generation of varied parallel scenarios.
@@ 2 Related work
2.1 Visual Generative Models
@@
2.2 Diffusion-Based Image Generation
@@
2.3 Remote Sensing Image Generation
@@
2.4 Remote Sensing Foundation Model
The foundation models have recently been preliminarily studied in the field of remote sensing, emerging as a new research direction. These models are per-trained on large amounts of data and can be adapted to various downstream tasks through fine-tuning, few-shot learning, and zero-shot learning. Current research on foundation models in remote sensing mainly focuses on two aspects: vision-centric foundation models and vision-language foundation models.
In earlier years, vision-centric foundation models in remote sensing are typically trained using supervised learning. However, these models are constrained by the limited availability of supervised training data. Recently, self-supervised techniques have become mainstream, as they can leverage large amounts of unlabeled data. The self-supervised pre-trained foundation models are designed to learn robust feature representations from unlabeled remote sensing images. They can then be transferred to various downstream tasks, such as image classification, object detection, and segmentation, resulting in improved performance. These models predominantly rely on the Masked Autoencoders (MAE) architecture[40] [41, 42, 43] and are tailored to exploit the distinctive features inherent in remote sensing imagery, including image content [44], multi-scale properties [45], temporal dynamics [46], and multispectral attributes [47, 42]. Besides, the improvement of the model’s structure and training paradigm [48] is used to enhance the model’s capacity for the representation of remote sensing data.
Remote sensing vision-language foundation models are trained on a large amount of image-text paired data. Some methods [49, 50, 51] adopt contrastive learning, which aims to train an image encoder and a text encoder to align the visual and linguistic domains. The learned aligned vision-language representations can be used for many downstream tasks, such as zero-shot classification [49], image-text retrieval [52, 53], and image captioning [54, 55]. For example, Zhang et al. [50] filtered remote sensing image-text pairs from publicly available large datasets to build the RS5M dataset and fine-tuned the CLIP [56] model to obtain a GeoRSCLIP model. Recently, autoregressive pre-training has been widely used for training vision-language models, such as [57, 58, 59, 60]. Most methods utilize CLIP to extract visual features and design connectors to map these visual features into large language models for predicting corresponding text. They typically start with image-text pre-training and then further refine the model through instruction fine-tuning.
Although the aforementioned work has made preliminary explorations in foundational models for remote sensing, research on generative foundational models in this field is still relatively scarce. In this paper, we propose MetaEarth, a novel generative foundation model for global-scale remote sensing image generation, which expands the research boundaries of current remote sensing foundation models.
3 METHODOLOGY
@@
3.1 Revisit of Denoising Diffusion Probabilistic Model
@@
3.2 The MetaEarth Self-Cascaded Generation Framework
@@
where the image $x_{0}^{(k+m)}$ generated from $(k+m)$-th stage will eventually has a size of $(N^{m}H\times N^{m}W)$ pixels. To achieve the above process with reasonable memory and computational costs, we designed a sliding window generation process and a noise sampling strategy that enables the generation of continuous, unbounded scenes in a memory-efficient manner. These approaches will be detailed in Subsection 3.3.
@@
3.3 Unbounded Image Generation
@@
3.4 Implimentation Details
3.4.1 Denoising Network Design
@@
For each feature map $h$ within every ResBlock, we divide the embedding $e_{t}^{(k)}$ into two parts using the chunk function in PyTorch and then scale the feature map $h$ with the chunked embeddings:
$\begin{split}&e_{\text{scale}},e_{\text{shift}}={\texttt{chunk}}(e_{t}^{(k)},2)\\ &h\leftarrow{(1+e_{\text{scale}})}\cdot h+e_{\text{shift}}\end{split}$ (22)
@@ 3.4.2 Training Details
4 EXPERIMENTS
4.1 Experimental Setup
4.1.1 Dataset
@@ 4.1.2 Evaluation Metrics
4.2 Qualitative Analysis
4.2.1 Global-Scale Image Generation
Our experiments show that the proposed MetaEarth can generate a variety of remote sensing scenes worldwide, including glaciers, snowfields, deserts, forests, beaches, farmlands, industrial areas, residential areas, etc. Fig. 1 and Fig. 5 showcase some results of our model. Owing to the guidance from the conditional input, the model is capable of generating vivid and diverse images with distinct regional characteristics at different resolutions. We also evaluated the semantical guidance provided by the image conditional input. In Fig. 6, we use large-scale low-resolution images of 256m/pixel in the Americas, Europe, and Asia as image condition guidance, and we can see the model can generate urban images consistent with the regional style.
@@ 4.2.1 Global-Scale Image Generation
4.2.2 Multi-Resolution Image Generation
@@
4.2.3 Arbitrary-sized Image Generation
@@
4.2.4 Generalization to Unseen Conditions
@@
4.2.5 Comparison with other text-to-image models
@@ 4.3 Ablation Study
4.3.1 Effectiveness of Resolution Guidance
@@
4.3.2 Evaluation on Noise Sampling
@@
4.4 Evaluation on Downstream Tasks